OpenStudy (amenah8):

Implicit differentiation: x^3 + y^3 = 3xy^2

OpenStudy (johnweldon1993):

Well just set it up \[\large \frac{d}{dx} (x^3 + y^3)= \frac{d}{dx} (3xy^2)\] \[\large \frac{d}{dx} x^3 + \frac{d}{dx} y^3= \frac{d}{dx} 3xy^2\] What would we do next?

OpenStudy (amenah8):

i'm not sure... :(

OpenStudy (amenah8):

do we have to get the y's to one side?

OpenStudy (johnweldon1993):

Well we have to differentiate first :) I know you can handle the first term...what is the derivative of x^3 with respect to x?

OpenStudy (amenah8):

3x^2 dx

OpenStudy (johnweldon1993):

So that's the first term! \[\large 3x^2 + \frac{d}{dx} y^3 = \frac{d}{dx} 3xy^2\] Now onto the second term \[\large \frac{d}{dx} y^3\] What is this derivative? Hint...use the chain rule

OpenStudy (johnweldon1993):

Let me know if you need help there :)

OpenStudy (amenah8):

3y^2 dy/dx?

OpenStudy (johnweldon1993):

Perfect!!! That's our second term...finish it up for me...what is the final term? \[\large \frac{d}{dx} 3xy^2\]

OpenStudy (amenah8):

3(2y) dy/dx ?

OpenStudy (johnweldon1993):

Not quite...we have to use the product rule there \[\large \frac{d}{dx} 3xy^2 = 3[y^2\frac{d}{dx}x + x\frac{d}{dx}y^2]\]

OpenStudy (amenah8):

3y^2 x + xy^2

OpenStudy (amenah8):

wait i meant: 3y^2 x + 3xy^2

OpenStudy (johnweldon1993):

Not quite *ignoring the 3 to make it look nicer* \[\large y^2 \frac{d}{dx} x + x \frac{d}{dx} y^2\] The derivative of x with respect to x is 1 \[\large y^2 + x\frac{d}{dx}y^2\] Using the chain rule for \(\large \frac{d}{dx}y^2\) what do we get?

OpenStudy (amenah8):

2y dy/dx

OpenStudy (johnweldon1993):

There we go So stay with me as I bring all the terms back \[\large 3x^2 + 3y^2 \frac{dy}{dx} = 3(y^2 + 2xy\frac{dy}{dx})\] Everything look good?

OpenStudy (amenah8):

i get it!

OpenStudy (johnweldon1993):

Perfect...now just solve for \(\large \frac{dy}{dx}\) and you're all set!

OpenStudy (amenah8):

could you please wait while i try to get the answer?

OpenStudy (johnweldon1993):

Not a problem, post back here when you're done :)

OpenStudy (amenah8):

could i simplify 3(y2+2xy dy/dx) to 3y^2 + 6xy dy/dx ?

OpenStudy (johnweldon1993):

Absolutely, that is what I would do!

OpenStudy (amenah8):

ok! just give me a moment

OpenStudy (amenah8):

dy/dx = (3xy^2 - 3x^2)/(3xy) ?

OpenStudy (johnweldon1993):

Hmm let's see here \[\large 3x^2 + 3y^2 \frac{dy}{dx} = 3y^2 + 6xy\frac{dy}{dx}\] Get the \(\large \frac{dy}{dx}\) to one side and the rest on the other \[\large 3x^2 - 3y^2 = 6xy\frac{dy}{dx} - 3y^2\frac{dy}{dx}\] Factor out the common term \[\large 3x^2 - 3y^2 = \frac{dy}{dx}(6xy - 3y^2)\] And finally divide both sides by \(\large 6xy - 3y^2\) \[\large \frac{dy}{dx} = \frac{3x^2 - 3y^2}{6xy - 3y^2}\] We can even simplify this farther by factoring out a 3 from top and bottom \[\large \frac{dy}{dx} = \frac{\cancel{3}(x^2 - y^2)}{\cancel{3}(2xy - y^2)}\] Giving us \[\large \frac{dy}{dx} = \frac{x^2 - y^2}{2xy - y^2}\]

OpenStudy (amenah8):

thank you so much! i see what i did wrong. i had subtracted the 3y^2 dy/dx wrong.

OpenStudy (johnweldon1993):

Not a problem! Glad it all made sense!

OpenStudy (amenah8):

thanks again!